Circular No. 5927

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD L.-F. Wang, Beijing Observatory; and E. J. Wampler, European South-

ern Observatory, report: “Observations of SN 1987A with the ESO 3.5-m New Technology Telescope in 1992 Jan. and in 1993 Jan. and Dec. indicate that, while the integrated brightness of the ring-shaped nebulosity surrounding SN 1987A has decreased by ∼ 20 percent during this 2-yr period, the distribution of light along the ring has changed dramatically from early observations. The ring is increasing in brightness at several locations. Following image deconvolution to achieve an effective resolution of 0′′.2, the 1993 Dec. frames show that the ring emission regions are now highly clumped and have a somewhat different structure when seen in Hα, [N II], or [O III]. Overall, the ring is brightening in the northwest, while continuing to fade in the south. The major new patches coincide roughly with the radio structure reported by Staveley-Smith et al. (1993, Nature 366, 136). These changes may herald the beginning of the interaction between the supernova ejecta and the nebular material surrounding SN 1987A. But significant amounts of SN 1987A envelope gas have not yet reached the ring, since spectra show that the nebular lines are still narrow. Perhaps the interactions between the expanding SN envelope and the residual gas interior to the ring that are producing the radio (Staveley-Smith et al. 1993) and x-ray (Beuermann et al. 1994, A.Ap. 281, L45) emissions also generate sufficient ultraviolet photons to reionize patches of gas in the ring. Near-infrared images now resolve the inner photosphere of SN 1987A. A comparison between the image profile of SN 1987A and those of field stars in deconvolved Gunn i frames with an effective resolution of 0′′.3 shows a resolved image with FWHM diameter 0′′.37.”

KR AURIGAE A. Antov and V. Popov, Department of Astronomy and National As-

tronomical Observatory, Bulgarian Academy of Sciences, Sofia, report that this object has faded from B = 13.8 in Dec. 1993 to B = 14.5 on 1994 Jan. 20, and that further fading to mag 17-18 can be expected. Spectroscopic and high-speed-photometric monitoring during the minimum state could be helpful in determining if the cessation of accretion is the reason for brightness dips.

1994 January 27 Daniel W. E. Green